OpenStudy (anonymous):

PLEASE HELPPP. I have no idea how to do this. log(10^3x^7) A.21log(x) B.3x^7 C.21 D.3+7log(x)

OpenStudy (whpalmer4):

remember that the log of a product is equal to the sum of the logs of the multiplicands

OpenStudy (whpalmer4):

@julian25 why did you delete? that was good stuff

OpenStudy (anonymous):

yes but i saw u telling me dont give the answer.........

OpenStudy (anonymous):

I honestly have no idea how to do this, or any idea with logs because I am taking an online program and have no base of trig/precal. First time taking it.): If you could please teach me.

jimthompson5910 (jim_thompson5910):

first you use the rule that log(A*B) = log(A) + log(B)

jimthompson5910 (jim_thompson5910):

so that will break log(10^3x^7) into log(10^3) + log(x^7)

jimthompson5910 (jim_thompson5910):

what is your next move?

OpenStudy (anonymous):

Do you just simply add them?

jimthompson5910 (jim_thompson5910):

what do you mean by that

OpenStudy (anonymous):

Wait is the answer 3x^7?

jimthompson5910 (jim_thompson5910):

what does log(10^3) simplify into

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

keep in mind that this log is by default base 10

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so log(10^3) + log(x^7) becomes 3 + log(x^7)

jimthompson5910 (jim_thompson5910):

how would you simplify log(x^7)

OpenStudy (anonymous):

I just put it into my calculator right now aha, and got 7.But how do you do it without a calculator?

jimthompson5910 (jim_thompson5910):

well you use the rule that log(x^y) = y*log(x)

jimthompson5910 (jim_thompson5910):

that will bring down the exponent so log(x^7) = 7*log(x)

jimthompson5910 (jim_thompson5910):

so in the end, the final answer is 3 + 7*log(x)

OpenStudy (anonymous):

I understand now! Thank you so much for the help! (:

jimthompson5910 (jim_thompson5910):

yw